OpenStudy (anonymous):

hep with this function, trying to find b f(x)=(1/3) b^x;f(1/2)=20

OpenStudy (campbell_st):

ok... so you know that \[20 = \frac{1}{3} \times b^{\frac{1}{2}}\] multiply both sides by 3 then square both sides to find the value of b

jimthompson5910 (jim_thompson5910):

I noticed this question was posted by two different people. Were you both able to figure this one out?

OpenStudy (anonymous):

whaaa? fo reals? who else posted it?

jimthompson5910 (jim_thompson5910):

I think you found it lol but were you able to figure it out?

OpenStudy (anonymous):

first tell me if it was a proper technique?

jimthompson5910 (jim_thompson5910):

yes it's the way to do it

jimthompson5910 (jim_thompson5910):

campbell_st is on the right track

OpenStudy (anonymous):

i got 3600

OpenStudy (anonymous):

would you say that is correct?

jimthompson5910 (jim_thompson5910):

f(x)=(1/3) b^x 20=(1/3) b^(1/2) 20*3 = b^(1/2) 60 = b^(1/2) sqrt(b) = 60 b = 60^2 b = 3600 so you got it

OpenStudy (anonymous):

thanks guys, i love it here. ya'll make learning fun!